OpenStudy (jaydenthequiet):

A car dealer has 7 new sedans she wants to display but room for only 5 on the showroom floor. How many ways can she display the 7 cars on the showroom floor? A. 5040 B. 2520 C. 35 D. 21 thanks

OpenStudy (mathmale):

This sounds like: "A car dealer has 7 new sedans she wants to display but room for only 5 on the showroom floor. How many combinations from 7 cars, 5 taken at a time, are there? We can write this mathematically as \[_{7}C _{5}\]

OpenStudy (mathmale):

How would you evaluate that? If you're not sure, look up "combinatorics" or "combinations and permutations" on the 'Net or in your online learning materials.

OpenStudy (jaydenthequiet):

8?

OpenStudy (mathmale):

How did you get 8? Are you familiar with the formula for calculating the number of combinations from n objects taken x at a time?

OpenStudy (jaydenthequiet):

I thought I was

OpenStudy (mathmale):

\[_{n}C _{x}=\frac{ n! }{ x!(n-x)! }\]

OpenStudy (mathmale):

Here, n = 7 and x = 5.

OpenStudy (jaydenthequiet):

\[\frac{ 7! }{ 5!(7-5)! }\]

OpenStudy (mathmale):

Good. Note that 7! = 7*6*5! and that this reduces to 7*6 if you divide it by that 5! in the denominator. That leaves 7*6 = 42 in the numerator; you still have to divide that by (7-5)!. What's that?

OpenStudy (mathmale):

In other words, what is the value of\[\frac{ 42 }{ 2! }?\]

OpenStudy (jaydenthequiet):

21

OpenStudy (mathmale):

that looks reasonable to me. What do YOU think?

OpenStudy (jaydenthequiet):

I just did it on my calculator and it says that twenty-one is correct

OpenStudy (mathmale):

Cool. Congrats! Have to get off the 'Net now. Hope to work with you again soon.